# Bilevel Reinforcement Learning via the Development of Hyper-gradient without Lower-Level Convexity

### v1

@@ 6 Theoretical Analysis
In this section, we prove global convergence and give the iteration complexity for the proposed algorithms. The analysis requires only the regularity conditions—specifically, boundedness and Lipschitz continuity—of the first-order information. This distinguishes from general AID-based methods, which require additional smoothness assumptions regarding second-order derivatives [23, 12, 39, 16, 38]. The properties of $\varphi$, $\pi^{*}(x)$, and $V^{*}(x)$ are outlined in Appendix F. Moreover, it is worth emphasizing that the analysis for M-SoBiRL is different from that for SoBiRL. Thus, we organize their properties and the detailed proofs in Appendix G and Appendix H.
@@
In the model-based scenario (4), $f$ is continuously differentiable. The gradient $\nabla f(x,\pi)$ is $L_{f}$-Lipschitz continuous, i.e., for any $x_{1},x_{2}\in\mathbb{R}^{n},\pi_{1},\pi_{2}\in\Delta^{\left|{\mathcal{S}}\right|}$,
$\left\|{\nabla f(x_{1},\pi_{1})-\nabla f(x_{2},\pi_{2})}\right\|\leq L_{f}\left({\left\|{x_{1}-x_{2}}\right\|_{2}+\left\|{\pi_{1}-\pi_{2}}\right\|_{\infty}}\right),$
and $\left\|{\nabla_{\pi}f(x,\pi^{*}(x)}\right\|_{2}\leq C_{f\pi}$.
@@
In the model-free scenario with upper-level function (12), $l(d_{1},d_{2},\ldots,d_{I};x)$ parameterized by $x$ is a function of $I$ trajectories, each consisting of $H$ steps. $l\left({d_{1},d_{2},\ldots,d_{I};x}\right)$ is bounded by $C_{l}$, i.e., $\left|{l\left({d_{1},d_{2},\ldots,d_{I};x}\right)}\right|\leq C_{l}$, and is continuously differentiable with respect to $x$. $l(d_{1},d_{2},\ldots,d_{I};x)$ is $L_{l}$-Lipschitz continuous, i.e., for any $x_{1},x_{2}\in\mathbb{R}^{n}$ and $\left({d_{1},d_{2},\ldots,d_{I}}\right)$,
$\left\|{l(d_{1},d_{2},\ldots,d_{I};x_{1})-l(d_{1},d_{2},\ldots,d_{I};x_{2})}\right\|_{2}\leq L_{l}\left\|{x_{1}-x_{2}}\right\|_{2},$
Moreover, $\nabla l(d_{1},d_{2},\ldots,d_{I};x)$ is $L_{l1}$-Lipschitz continuous, i.e., for any $x_{1},x_{2}\in\mathbb{R}^{n}$ and $\left({d_{1},d_{2},\ldots,d_{I}}\right)$,
$\left\|{\nabla l(d_{1},d_{2},\ldots,d_{I};x_{1})-\nabla l(d_{1},d_{2},\ldots,d_{I};x_{2})}\right\|_{2}\leq L_{l1}\left\|{x_{1}-x_{2}}\right\|_{2}.$
Assumptions 6.1 and 6.2 specify the requirements for the upper-level problems in model-based and model-free scenarios, respectively. Generally, the boundedness and Lipschitz continuity assumptions related to $f$ are standard in bilevel optimization [23, 37, 3, 13, 45].
@@
For any $x\in\mathbb{R}^{n},\ (s,a)\in\mathcal{S}\times\mathcal{A}$, the reward is bounded by $C_{r}$, i.e., $\left|{r_{sa}(x)}\right|\leq C_{r}$, and the gradient is bounded by $C_{rx}$, i.e., $\left\|{\nabla r_{sa}(x)}\right\|_{2}\leq C_{rx}$. Additionally, $r_{sa}(x)$ is $L_{r}$-Lipschitz smooth, i.e., for any $x_{1},x_{2}\in\mathbb{R}^{n}$, $\left\|{\nabla r_{sa}(x_{1})-\nabla r_{sa}(x_{2})}\right\|_{2}\leq L_{r}\left\|{x_{1}-x_{2}}\right\|_{2}$.
This assumption on the parameterized reward function is common in RL [85, 1, 10, 69].
@@
Under Assumptions 6.1 and 6.3, in Algorithm 1, we can choose constant step sizes $\beta,\xi$, and the inner iteration number $N\sim\mathcal{O}(1)$. Then the iterates $\{x_{k}\}$ satisfy
@@
Regarding the model-free scenario, coupled with Proposition 5.2, the next proposition marks a substantial development to clarify the Lipschitz property of the hyper-gradient.
@@
$\left\|\nabla Q_{sa}^{*}(x_{1})-\nabla Q_{sa}^{*}(x_{2})\right\|_{2}\leq L_{V1}\left\|x_{1}-x_{2}\right\|_{2},\quad\left\|\nabla V_{s}^{*}(x_{1})-\nabla V_{s}^{*}(x_{2})\right\|_{2}\leq L_{V1}\left\|x_{1}-x_{2}\right\|_{2},$
and
with $L_{V1}$ and $L_{\phi}$ specified in Propositions H.5 and H.6, respectively.
@@
Under Assumptions 6.2 and 6.3 and given the accuracy $\epsilon$ in Algorithm 2, we can set the constant step size $\beta<\frac{1}{2L_{\phi}}$, then the iterates $\{x_{k}\}$ satisfy
$\frac{1}{K}\sum_{k=1}^{K}{\left\|\nabla\phi\left(x_{k}\right)\right\|^{2}}\leq\frac{\phi\left(x_{1}\right)-\phi^{*}}{K\left(\frac{\beta}{2}-\beta^{2}L_{\phi}\right)}+\frac{1+2\beta L_{\phi}}{1-2\beta L_{\phi}}L^{2}_{\widetilde{\phi}}\ \epsilon\ ,$
@@
Combined with Table 1, the convergence property in Theorem 6.6 implies that the proposed algorithm SoBiRL realizes better iteration complexity than PBRL [68] with the same computation cost at each outer and inner iteration. Moreover, SoBiRL attains the convergence rate of the same order as PARL [10], but only employs first-order oracles and gets rid of additional assumptions on the lower-level problem.
7 Experiment
In this section, we conduct an experiment on the Atari game, BeamRider from the Arcade Learning Environment (ALE) [5] to empirically validate the efficiency of the model-free algorithm, SoBiRL. The reward provided by ALE serves as the ground truth, and for each trajectory pair, preference is assigned to the trajectory with higher accumulated ground-truth reward. We compare SoBiRL with DRLHF [14] and a baseline algorithm SAC [29, 30]. Both SoBiRL and DRLHF adopt deep neural networks to predict rewards, while the baseline SAC receives ground-truth rewards for training. More experiment details and parameter settings are summarized in Appendix I.
Figure 1: Comparison of algorithms on the Atari game, BeamRider, evaluated by the ground-truth reward. Both SoBiRL and DRLHF collect a total of $3000$ trajectory pairs. The running average over $15$ consecutive episodes is adopted for the presentation, and results are averaged over $5$ seeds.
@@ 7 Experiment
The results are presented in Figure 1, showing that even in the face of unknown rewards, the performance of SoBiRL is on a par with the baseline SAC. Additionally, SoBiRL achieves a higher episodic return and a longer episodic life than DRLHF within the same outer timesteps. Another interesting observation is that, compared to the ground truth, the preference prediction based on the reward model of DRLHF achieves an accuracy of approximately $85\%$, while it hovers around $54\%$ with the reward model of SoBiRL. This stems from different update rules. Specifically, DRLHF alternates between learning a policy given the parameterized reward $r(x)$ and minimizing $\mathbb{E}\left[{l_{h}}\right]$ based on collecting trajectories, which decouples the task into two separate phases. It aligns preferences with the ground truth, but overlooks the exploration for rewards promoting better policy. Nevertheless, SoBiRL glues the two levels via the hyper-gradient ${\nabla}{\phi}$, with the first term (13) exploiting the preference information to align the reward predictor and the second term (14) exploring the implicit reward to unearth a more favorable policy. Therefore, SoBiRL produces superior results in reward prediction and policy optimization, although it may exhibit lower accuracy in alignment of trajectory preference.
Appendix
@@ Appendix A Related Work in Bilevel Optimization
By the implicit function theorem, the approximate implicit differentiation (AID) based method regards the optimal lower-level solution as a function of the upper-level variable, which yields the hyper-gradient to instruct the upper-level update. It implements alternating (inexact) gradient descent steps between two levels [23, 37, 13, 16, 45, 32]. In addition, generalizing the lower-level strong convexity, the studies [24, 25, 26] focused on the bilevel problem where the lower-level problem is formulated as a fixed-point equation. Moreover, a line of research has been dedicated to the bilevel problem with a nonconvex lower-level objective [51, 68, 35, 56, 53]. Specifically, Shen and Chen [68] proposed a penalty-based method to bypass the requirement of lower-level convexity. In [35], a momentum-based approach was designed to solve the bilevel problem with a lower-level function satisfying the PL condition. Liu et al. [53] utilized the Moreau envelope based reformulation to provide a single-loop algorithm for nonconvex-nonconvex bilevel optimization.

### v2

@@ 5 MODEL-FREE SOFT BILEVEL REINFORCEMENT LEARNING
6 THE STOCHASTIC EXTENSION
The preceding sections establish bilevel methods through the lens of deterministic optimization. However, in practical implementations of reinforcement learning, quantities are typically approximated via sampling data, e.g., ~$\widetilde{\nabla}\phi(x_{k},\pi_{k})$ defined in an expected form. As a result, the sampling process introduces stochasticity to the algorithm. Therefore, to enrich the developed framework SoBiRL both practically and theoretically, it is essential to extend it to stochastic settings. In detail, we approach this in two steps. (1) Designing a scheme to estimate the hyper-gradient $\nabla\phi$ by sampling trajectories and then characterizing the resulting stochasticity, i.e., the bias and variance. (2) Replacing the hyper-gradient in SoBiRL with its stochastic counterpart and incorporating a momentum technique to accommodate the stochasticity, leading to our stochastic variant, Stoc-SoBiRL.
In the stochastic scenario, we still focus on the problem formulation (4) with the upper-level function (11) as discussed in Section 5. Denote the trajectory tuple $(d_{1},d_{2},\ldots,d_{I})$ as $\bm{d}$ for simplicity. In the $k$-th outer iteration, to estimate $\nabla\phi$, an expectation with respect to the random variable $\bm{d}$, we can generate $M$ independent tuples $\bm{d}^{m}:=(d^{m}_{1},d^{m}_{2},\ldots,d^{m}_{I})$ for $m=1,2,\ldots,M$, evaluate corresponding quantities along each $\bm{d}^{m}$ and average them. To this end, it is necessary to tackle the terms $\nabla Q^{*}$ and $\nabla V^{*}$ in (13). Assuming access to a generative model (Kearns and Singh,, 1998; Li et al., 2024a, ), from any initial state-action pair $(s,a)$, we sample $J$ independent trajectories of length $T$ by implementing $\pi_{k}$, i.e., for $j=1,2,\ldots,J$,
$\xi_{k}^{j}(s,a):=\{(s_{0}^{j}=s,a_{0}^{j}=a);(s_{1}^{j},a_{1}^{j}),\ldots,(s_{T-1}^{j},a_{T-1}^{j})\}.$
Collecting all these random variables yields $\bm{\xi}_{k}:=\{\xi_{k}^{j}\left(s,a\right),j=1,\ldots,J,s\in\mathcal{S},a\in\mathcal{A}\}$, and the estimator for $\nabla Q^{*}$ can be constructed as follows.
¯$\bar{\nabla}Q_{sa}^{\bm{\xi}_{k}}\left(x_{k},\pi_{k}\right):=\frac{1}{J}\sum_{j=1}^{J}{\sum_{t=0}^{T-1}{\gamma^{t}\nabla r_{s_{t}^{j}a_{t}^{j}}\left(x_{k}\right)}}.$
The random variable $\bm{\zeta}_{k}$ and the associated estimator ¯$\bar{\nabla}V^{\bm{\zeta}_{k}}$ for $\nabla V^{*}$ are constructed similarly. Consequently, denoting $\bm{D}_{k}:=(\bm{d}_{k}^{1},\bm{d}_{k}^{2},\ldots,\bm{d}_{k}^{M})$, we obtain the stochastic hyper-gradient:
¯$\bar{\nabla}\phi\left(\bm{D}_{k},\bm{\xi}_{k},\bm{\zeta}_{k};x_{k},\pi_{k}\right)=\frac{1}{M}\sum_{m=1}^{M}{\nabla l\left(\bm{d}_{k}^{m};x_{k}\right)}+\frac{1}{\tau M}$
×$\times\!\sum_{m=1}^{M}{l\left(\bm{d}^{m}_{k};x_{k}\right)}\sum_{i}\sum_{h}{\bar{\nabla}\!\left(Q_{s_{h}^{m,i}a_{h}^{m,i}}^{\bm{\xi}_{k}}-V_{s_{h}^{m,i}}^{\bm{\zeta}_{k}}\right)}(x_{k},\pi_{k}),$
which is abbreviated as ¯$\bar{\nabla}\phi_{k}$ in the following discussion. The sampling scheme is outlined in Algorithm 3.
Momentum techniques are known to be beneficial for reducing variance and accelerating algorithms (Cutkosky and Orabona,, 2019). To this end, we maintain a momentum-instructed $h_{k}$ in the $k$-th outer iteration,
$h_{k}\!=\!\bar{\nabla}\phi_{k}+(1-\mu_{k})(h_{k-1}-\bar{\nabla}\phi(\bm{D}_{k},\bm{\xi}_{k},\bm{\zeta}_{k};x_{k-1},\pi_{k-1})),$
which tracks ¯$\bar{\nabla}\phi_{k}$ via current and historical hyper-gradient estimates. Using $h_{k}$ to update the upper-level $x_{k}$, we obtain the stochastic Algorithm 4, Stoc-SoBiRL.
Figure 1: Comparison of algorithms on the Atari game, BeamRider, evaluated by the ground-truth reward. Each bilevel algorithm collects a total of $3000$ trajectory pairs. The running average over $15$ consecutive episodes is adopted for the presentation, and results are averaged over $5$ seeds.
@@ 7 THEORETICAL ANALYSIS
In this section, we prove global convergence and give the iteration and sample complexity for the proposed algorithms. The analysis requires only the regularity conditions—specifically, boundedness and Lipschitz continuity—of the first-order information. This distinguishes from general AID-based methods, which require additional smoothness assumptions regarding second-order derivatives (Ghadimi and Wang,, 2018; Chen et al.,, 2021; Khanduri et al.,, 2021; Dagréou et al.,, 2022; Ji et al.,, 2022). The properties of $\varphi$, $\pi^{*}(x)$, and $V^{*}(x)$ are outlined in Appendix F. Moreover, it is worth emphasizing that the analysis for M-SoBiRL is different from that for SoBiRL. Thus, we organize their properties and the detailed proofs in Appendix G and Appendix H. The analysis for Stoc-SoBiRL is referred to Appendix I.
@@
In the model-based scenario (4), $f$ is continuously differentiable. The gradient $\nabla f(x,\pi)$ is $L_{f}$-Lipschitz continuous, i.e., for any $x_{1},x_{2}\in\mathbb{R}^{n},\pi_{1},\pi_{2}\in\Delta^{\left|{\mathcal{S}}\right|}$, $\left\|{\nabla f(x_{1},\pi_{1})-\nabla f(x_{2},\pi_{2})}\right\|_{2}\leq L_{f}\left({\left\|{x_{1}-x_{2}}\right\|_{2}+\left\|{\pi_{1}-\pi_{2}}\right\|_{\infty}}\right)$. Additionally, the boundedness condition holds, $\left\|{\nabla_{\pi}f(x,\pi^{*}(x))}\right\|_{2}\!\leq\!C_{f\pi}$.
@@
In the model-free scenario with upper-level function (11), denote $\bm{d}=(d_{1},d_{2},\ldots,d_{I})$. $l(\bm{d};x)$ parameterized by $x$ is a function of $I$ trajectories, each consisting of $H$ steps. It is bounded, i.e., $\left|{l\left({\bm{d};x}\right)}\right|\leq C_{l}$, and is continuously differentiable with respect to $x$. $l(\bm{d};x)$ is $L_{l}$-Lipschitz continuous, i.e., for any $x_{1},x_{2}\in\mathbb{R}^{n}$ and $\bm{d}$, $\left\|{l(\bm{d};x_{1})-l(\bm{d};x_{2})}\right\|_{2}\leq L_{l}\|x_{1}-x_{2}\|_{2}.$ Moreover, $\nabla l(\bm{d};x)$ is $L_{l1}$-Lipschitz continuous, i.e., for any $x_{1},x_{2}\in\mathbb{R}^{n}$ and $\bm{d}$, $\left\|{\nabla l(\bm{d};x_{1})-\nabla l(\bm{d};x_{2})}\right\|_{2}\leq L_{l1}\left\|{x_{1}\!-\!x_{2}}\right\|_{2}$.
Assumptions 7.1 and 7.2 specify the requirements for the upper-level problems in model-based and model-free scenarios, respectively. Generally, the boundedness and Lipschitz continuity assumptions related to $f$ are standard in bilevel optimization (Ghadimi and Wang,, 2018; Ji et al.,, 2021; Arbel and Mairal,, 2022; Chen et al., 2022b, ; Li et al.,, 2022).
@@
The reward is bounded by $C_{r}$, i.e., $\left|{r_{sa}(x)}\right|\leq C_{r}$, and the gradient is bounded by $C_{rx}$, i.e., $\left\|{\nabla r_{sa}(x)}\right\|_{2}\leq C_{rx}$. Additionally, $r_{sa}(x)$ is $L_{r}$-Lipschitz smooth, i.e., for any $x_{1},x_{2}\in\mathbb{R}^{n}$ and $(s,a)\in\mathcal{S}\times\mathcal{A}$, $\left\|{\nabla r_{sa}(x_{1})-\nabla r_{sa}(x_{2})}\right\|_{2}\leq L_{r}\left\|{x_{1}-x_{2}}\right\|_{2}$.
This assumption on the parameterized reward function is common in RL (Zhang et al.,, 2020; Lan et al.,, 2022; Chakraborty et al.,, 2024; Shen et al.,, 2024).
@@
Under Assumptions 7.1 and 7.3, in M-SoBiRL, we can choose constant step sizes $\beta,\eta$, and the inner iteration number $N\sim\mathcal{O}(1)$. Then the iterates $\{x_{k}\}$ satisfy
$\frac{1}{K}\sum_{k=1}^{K}\left\|{\nabla\phi(x_{k})}\right\|_{2}^{2}=\mathcal{O}\left({K^{-1}}\right).$
@@
Coupled with Proposition 5.2, the next proposition marks a substantial development to clarify the Lipschitz property of the hyper-gradient.
@@
Under Assumption 7.3, we have $\left\|{\nabla\phi(x_{1})-\nabla\phi(x_{2})}\right\|_{2}\leq L_{\phi}\left\|{x_{1}-x_{2}}\right\|_{2}$ for any $x_{1},x_{2}\in\mathbb{R}^{n}$, with $L_{\phi}$ specified in Proposition H.6.
@@
Under Assumptions 7.2 and 7.3 and given the accuracy $\epsilon$ in Algorithm 1, we can set the constant $\beta<\frac{1}{2L_{\phi}}$, and then the iterates $\{x_{k}\}$ satisfy
$\frac{1}{K}\sum_{k=1}^{K}{\left\|\nabla\phi\left(x_{k}\right)\right\|_{2}^{2}}\leq\frac{\phi\left(x_{1}\right)-\phi^{*}}{K\left(\frac{\beta}{2}-\beta^{2}L_{\phi}\right)}+\frac{1+2\beta L_{\phi}}{1-2\beta L_{\phi}}L^{2}_{\widetilde{\phi}}\ \epsilon\,,$
@@
Combined with Table 1, the convergence property in Theorem 7.6 implies that SoBiRL realizes better iteration complexity than PBRL (Shen and Chen,, 2023) with the same computation cost at each outer and inner iteration. Moreover, SoBiRL attains the convergence rate of the same order as PARL (Chakraborty et al.,, 2024), but only employs first-order oracles and gets rid of the convexity assumption on the lower-level problem.
In the context of stochastic bilevel RL, the distribution of samples $(\bm{D}_{k},\bm{\xi}_{k},\bm{\zeta}_{k})$ relies on the variable $\pi_{k}$, which differs from related work employing momentum techniques (Cutkosky and Orabona,, 2019; Khanduri et al.,, 2021; Huang,, 2023). To circumvent this misalignment in Stoc-SoBiRL, we give the statistical properties (bias and variance of the stochastic hyper-gradient) in Section I.2 and the convergence analysis in Section I.3.
Theorem 7.7 (Stochastic).
Under Assumptions 7.2 and 7.3, we can choose appropriate sampling configurations $M\sim\mathcal{O}(K^{4/3}),\,J\sim\mathcal{O}(1),\,T\sim\mathcal{O}(\log K)$. Then the iterates $\{x_{k}\}$ generated by Algorithm 4 satisfy
$\frac{1}{K}\sum_{k=1}^{K}{\mathbb{E}\left[\left\|\nabla\phi(x_{k})\right\|_{2}^{2}\right]}=\mathcal{O}\left(\frac{\log K}{K^{2/3}}\right)=\widetilde{\mathcal{O}}\left(K^{-\frac{2}{3}}\right).$
Detailed parameter setting is given in Theorem I.7.
Enlightened by the momentum acceleration, the average norm square of the hyper-gradient achieves $\epsilon$ within $\widetilde{\mathcal{O}}(\epsilon^{-1.5})$ outer iterations. Therefore, the samples required by the upper-level leader are of the order $\widetilde{\mathcal{O}}(\epsilon^{-3.5})$, which can be alleviated in practice, as the leader can leverage vectorized environments (Makoviychuk et al.,, 2021) within each outer iteration; see Appendix L for further discussion.
8 EXPERIMENTS
In this section, we conduct RLHF experiments on Atari games and Mujoco environments to validate the efficiency of SoBiRL, and one synthetic experiment to verify the convergence of M-SoBiRL. The experiment details and parameter settings are provided in Appendix J. We have made the code available on https://github.com/UCAS-YanYang/SoBiRL.
Table 2: Comparison of algorithms on the Atari games and Mujoco simulations, evaluated by the ground-truth reward. Each bilevel algorithm collects a total of 3000 trajectory pairs. The results are recorded after appropriate timesteps and averaged over 5 seeds.
Algorithm BeamRider (4M ) Seaquest (1M ) SpaceInvaders (1M ) HalfCheetah (1M ) Walker2d (1M ) Hopper (500k )
SoBiRL $8459.7\pm 1989.1$ $746.7\pm 16.4$ $595.6\pm 23.8$ $4283.6\pm 363.0$ $2996.0\pm\phantom{0}97.4$ $2812.4\pm 217.5$
DRLHF $6124.5\pm\phantom{0}433.1$ $561.9\pm 38.0$ $483.9\pm 27.3$ $3466.7\pm 424.2$ $2389.4\pm 101.8$ $2455.8\pm 457.2$
PBRL $6798.2\pm\phantom{0}784.0$ $651.2\pm 24.6$ $554.0\pm 33.4$ $4022.8\pm 311.8$ $2257.8\pm 127.9$ $2799.1\pm 155.5$
HPGD $7503.9\pm 1863.6$ $728.1\pm 57.5$ $601.8\pm 36.0$ $3975.8\pm 476.1$ $2487.5\pm 223.8$ $2803.3\pm 354.5$
SAC $9780.0\pm 1877.4$ $891.1\pm 31.3$ $783.8\pm 56.7$ $5769.9\pm 765.3$ $4552.5\pm\phantom{0}23.3$ $2207.6\pm\phantom{0}18.4$
@@ 8 EXPERIMENTS
In RLHF experiments, we compare SoBiRL with the bilevel algorithms DRLHF (Christiano et al.,, 2017), PBRL (Shen et al.,, 2024), HPGD Thoma et al., (2024), and a baseline algorithm SAC (Haarnoja et al., 2018a, ; Haarnoja et al., 2018b, ). All the bilevel solvers harness deep neural networks to predict rewards, while the baseline SAC receives ground-truth rewards for training. The results are reported in Figures 1, 2, 3, and Table 2, showing that even in the face of unknown rewards, the performance of SoBiRL is on a par with the baseline SAC. Additionally, SoBiRL achieves a higher episodic return than the compared methods within the same timesteps. Another interesting observation is that, taking BeamRider for example, compared to the ground truth, the preference prediction based on the reward model of DRLHF achieves an accuracy of approximately $85\%$, while it hovers around $54\%$ with the reward model of SoBiRL. This stems from different update rules. Specifically, DRLHF alternates between learning a policy given the parameterized reward $r(x)$ and minimizing $\mathbb{E}\left[{l_{h}}\right]$ based on collecting trajectories, which decouples the task into two separate phases. It aligns preferences, but overlooks the exploration for rewards promoting better policy. Nevertheless, SoBiRL glues the two levels via the hyper-gradient ${\nabla}{\phi}$, with the first term (12) exploiting the preference information to align the reward predictor and the second term (13) exploring the implicit reward to unearth a more favorable policy. Therefore, SoBiRL produces superior results in reward prediction, although it exhibits lower accuracy in alignment of trajectory preference.
Regarding M-SoBiRL, the results of a synthetic experiment are shown in Figure 4. The curves exhibit the benign convergence property of the proposed algorithm.
Acknowledgments
Bin Gao was supported by the Young Elite Scientist Sponsorship Program by CAST. Ya-xiang Yuan was supported by the National Natural Science Foundation of China (grant No. 12288201). The authors are grateful to the Program and Area Chairs and Reviewers for their detailed and valuable comments and suggestions.
Supplementary Materials
@@ Appendix A RELATED WORK IN BILEVEL OPTIMIZATION
By the implicit function theorem, the approximate implicit differentiation (AID) based method regards the optimal lower-level solution as a function of the upper-level variable, which yields the hyper-gradient to instruct the upper-level update. It implements alternating (inexact) gradient descent steps between two levels (Ghadimi and Wang,, 2018; Ji et al.,, 2021; Chen et al., 2022b, ; Dagréou et al.,, 2022; Li et al.,, 2022; Hong et al.,, 2023). In addition, generalizing the lower-level strong convexity, the studies (Grazzi et al.,, 2020, 2021, 2023) focused on the bilevel problem where the lower-level problem is formulated as a fixed-point equation. Moreover, a line of research has been dedicated to the bilevel problem with a nonconvex lower-level objective (Liu et al., 2021c, ; Shen and Chen,, 2023; Huang,, 2023; Lu,, 2023; Liu et al., 2024a, ). Specifically, Shen and Chen, (2023) proposed a penalty-based method to bypass the requirement of lower-level convexity. In (Huang,, 2023), a momentum-based approach was designed to solve the bilevel problem with a lower-level function satisfying the PL condition. Liu et al., 2024a  utilized the Moreau envelope based reformulation to provide a single-loop algorithm for nonconvex-nonconvex bilevel optimization. In stochastic settings, bilevel algorithms have been developed in (Ghadimi and Wang,, 2018; Ji et al.,, 2021; Khanduri et al.,, 2021; Chen et al., 2022b, ).